OpenStudy (livielaughs):

The probability of a basketball player hitting a foul shot is 2/3. How many shots would you expect her to make in 120 attempts? (the reason im asking dumb questions is because I am 100% LEFT brain :P)

OpenStudy (kiwibyrd):

your equation would be 2/3(120)

OpenStudy (kiwibyrd):

120/3= 40 40(2) = 80

OpenStudy (kiwibyrd):

80/120 = 2/3

OpenStudy (kiwibyrd):

80 is the number of foul shots

OpenStudy (livielaughs):

sorry some one came to the door to drop off food :P

OpenStudy (livielaughs):

and thankyou